How much copper and how much magnesium should be added to 200lbs of an alloy containing 20% copper and 40% magnesium to produce an alloy containing 24% copper and 50% magnesium?
Answer by stanbon(75887)

How much copper and how much magnesium should be added to 200lbs of an alloy containing 20% copper and 40% magnesium to produce an alloy containing 24% copper and 50% magnesium?
--------------------
Copper Equation:: 0.20*200 + 1*c = 0.24(200+c)
20*200 + 100c = 24*200 + 24c
76c = 4*200
c = 200/19
c = 10.53 lbs (amt. of pure copper to add)
================================================
Magnesium Eq::::: 0.40*200 + 1*m = 0.50(200+m)
40*200 + 100m = 50*200 + 50m
50m = 10*200
m = 40 lbs (amt of pure magnesium to add)
================================================
Cheers,
Stan H.
===============